OpenStudy (anonymous):

Hi All, is there anyone who might be able to show me how to check for accuracy on this problem I worked? 2y/3-3/4=5/12. The answer is y=7/4 so I do know I need to replace y with 7/4 and work it but I can't seem to figure out how to lay that out? Anyone? Really appreciate any help:) Denise

OpenStudy (unklerhaukus):

\[\frac{2y}{3}-\frac{3}{4}=\frac{5}{12}\]\[ \qquad\downarrow \]\[\frac{2\times(7/4)}{3}-\frac{3}{4}=\frac{5}{12}\]\[\frac{14}{12}-\frac{3}{4}=\frac{5}{12}\]\[\frac{14}{12}-\frac{9}{12}=\frac{5}{12}\]\[14-9=5\] which is true. therefore y = 7/4 is a solution

OpenStudy (anonymous):

oh thanks, sorry, got stuck on the phone, thank you soooooooooooo much, Denise:)

OpenStudy (unklerhaukus):

do you follow the method?

OpenStudy (anonymous):

no, how did you get 14/12 UR:)?

OpenStudy (anonymous):

I understand the LCD;) but don't see how you got the 14/12:)

OpenStudy (anonymous):

and we just cancel out the 12's as denominators, I get that part too:)

OpenStudy (unklerhaukus):

\[\frac{2\times 7/4}{3}\]\[=\frac{14/3}{4}\]\[=\frac{14}{3}\frac{1}{4}\]\[=\frac{14}{12}\]

OpenStudy (unklerhaukus):

generalizing \[\frac{a/b}{c/d}=\frac{ad}{bc}\]

OpenStudy (anonymous):

Oh, that is something I totally forgot how to do, I see what you did though. Can't say I could do another one but I'll try another and see;) practice practice right;)

OpenStudy (anonymous):

Thank you again, I think that was the last one I had to check and my homework is done:) Denise:)

OpenStudy (unklerhaukus):

simplify \[\frac{1}{3/b}\]

OpenStudy (anonymous):

sec, let me think a minute:)

OpenStudy (anonymous):

1/3 first right?

OpenStudy (unklerhaukus):

the simplified form is a single fraction

OpenStudy (unklerhaukus):

where has the b gone?

OpenStudy (unklerhaukus):

you can use the generalized formula above

OpenStudy (anonymous):

1/3 over b, nope, Im not getting it.

OpenStudy (anonymous):

right, I have it written here on my paper, sec

OpenStudy (anonymous):

arrrrrrrrrrrrrg:(

OpenStudy (unklerhaukus):

ok so the step is to multiply both the numerator and denominator by the same thing,

OpenStudy (anonymous):

you have to eliminate the b right?

OpenStudy (unklerhaukus):

\[\frac{1}{3/b}\] times both top and bottom by b , what do you get?

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

1b/3b and cancel out b's

OpenStudy (unklerhaukus):

\[ \frac{1}{3/b}\]\[ =1\times\frac{1}{3/b}\]\[ =\frac{b}{b}\times\frac{1}{3/b}\]\[=...\]

OpenStudy (unklerhaukus):

the simplified form will have a b in it

OpenStudy (anonymous):

oh:( Thought I had it, sec

OpenStudy (anonymous):

B1/3?

OpenStudy (unklerhaukus):

oh that is looking much better

OpenStudy (unklerhaukus):

remember\[1\times b=b\],so you can simplify one bit more

OpenStudy (unklerhaukus):

\[\frac{1}{3/b}=\frac{b\times1}{3}=...\]

OpenStudy (anonymous):

so it is just B/3!!

OpenStudy (unklerhaukus):

!!!!yes

OpenStudy (anonymous):

LOL:) Thank you, you are so kind to teach me that:) That will come in handy I can tell you, thank you so much:)

OpenStudy (unklerhaukus):

we had a fraction in a fraction, and we pulled the denominator in the denominator, into the numerator

OpenStudy (anonymous):

Ok, well, I better get finished up here. I will be back tomorrow night I suspect. You are not in Texas I hope?? Storms/Tornados?

OpenStudy (unklerhaukus):

hmm, no. i live in Australia

OpenStudy (anonymous):

Oh that is something, I can tell folks I got help all the way from Australia on my Algebra!! Nite UR, hope to see you again;) Denise

OpenStudy (unklerhaukus):

Until next time, all the best

OpenStudy (anonymous):

same to you:)